OpenStudy (anonymous):

there are 10 light bulbs in a bag. two are defective select a light bulb whhat is the probability of selecting a light bulb that is not defective

OpenStudy (amistre64):

1 - 2/10 = .8 ?

OpenStudy (anonymous):

it comes out as a fraction

OpenStudy (amistre64):

8/10 = 4/5 :)

OpenStudy (asadkarim7):

8/10

OpenStudy (anonymous):

dang your smart

OpenStudy (amistre64):

well, my other option was be cute; and well ..... you see which one i picked lol

OpenStudy (anonymous):

i got another problem for you i bet you dnt get this one